# ![[Uncaptioned image]](extracted/5527354/head_foot/header_bar.png) CryinGAN: Design and evaluation of point-cloud-based generative adversarial networks using disordered materials – application to LiScCl-LiCoO battery interfaces | Adrian Xiao Bin Yong, and Elif Ertekin |

### v1

@@ 4.1 GAN and Dataset
The dataset used for training the GANs consists of 1,500 relaxed low-interface-energy structures. All structures had the same lattice parameters and composition, so each structure was represented by its 2D matrix of fractional coordinates only. CryinGAN models were trained with = 0, 0.01, 0.05, and 0.1 for 100,000 epochs to study the effect of on model performance. Three separate models were trained for each value. A batch size of 32 was used, and Adam optimizers with learning rate of were used for the generator and discriminator(s). The generator was trained only once every 5 batches to help stabilize the training. CryinGAN-comb, CryinGAN-max, and CryinGAN-mix models were trained using the same procedure, and three separate models were trained for each model configuration. CryinGAN-max and CryinGAN-mix, which use two discriminators, were trained using = 0 and 0.05. For = 0, max pooling and mix pooling were used for the fractional coordinate discriminator of CryinGAN-max and CryinGAN-mix respectively. For = 0.05, max pooling and mix pooling were used for the bond distance discriminator of CryinGAN-max and CryinGAN-mix respectively, whereas average pooling was used for the fractional coordinate discriminator of both models. For each trained model, structures were generated, relaxed using the M3GNet interatomic potential 48, and the (normalized) interface energy distribution was calculated using 1,000 relaxed structures.
@@
We also tested a graph convolutional neural network architecture by adapting CGCNN 54 as the discriminator. Instead of predicting material properties, the output of CGCNN was used to estimate the Wasserstein distance between the training and generated structures. Default values for the hyperparameters were used, as provided in the code repository of CGCNN. The GAN was optimized using the Adam optimizer and a batch size of 32, where we tested learning rates of , , , and . We found that the losses diverged and the GAN was unable to generate meaningful structures. We did not further pursue the development of graph convolutional neural networks as discriminators, considering similar failures reported in literature for point cloud GANs 53.
@@
The best GAN architecture was found to be CryinGAN using both discriminators, where was found to be optimum among the values tested. We trained CryinGAN with to use for further structural analysis of the generated structures. The generator was trained once every 2 batches to speed up the training. We found the training to be stable at this frequency as shown by the Wasserstein distance plot in ESI Fig. S14a. Although the Wasserstein distance plateaued early in the training, the percentage of (relaxed) low-interface-energy structures continued to gradually increase (see ESI Fig. S14b and c). We stopped the training when the energy improvements have mostly plateaued, and used the last trained model as the best GAN model. Structures were generated using the model, and relaxed using M3GNet followed by DFT calculations to determine the interface energy distribution and perform subsequent structural analysis.
4.2 Structural analysis
Three datasets were used to perform the structural analysis: (1) CryinGAN-generated structures with low interface energy ( J/m), (2) randomly generated structures with low interface energy (i.e., the training structures), and (3) randomly generated structures with high interface energy ( J/m). Each dataset contains 1,500 structures that were relaxed using the M3GNet interatomic potential followed by DFT calculations.
@@ 4.2 Structural analysis
To evaluate the structural similarity between the CryinGAN dataset and the training dataset, we compared the coordination environment of atoms in the interface region. The same comparison was carried out between the high energy dataset and the training dataset. For each structure, the CrystalNNFingerprint 56 was calculated for Li and Sc atoms in the interface region, allowing only the anions (Cl and O) to be considered as neighbors. The CrystalNNFingerprint is a vector containing coordination likelihoods, and local structure order parameters (LoStOPs) that describe the extent to which an observed coordination environment matches a perfect coordination motif. For each atom type (Li/Sc), an average fingerprint was calculated by averaging the site fingerprints of each structure, then averaging across all structures in a dataset. The Euclidean distance and cosine similarity between the average fingerprint of the CryinGAN/high energy dataset and the training dataset were calculated. To visualize the coordination motif distribution of the datasets, the most likely coordination motif of each interface Li/Sc atom was identified by first selecting the coordination number with the highest likelihood, then selecting the coordination motif of this coordination number with the highest LoStOP. We only plotted coordination motifs that appeared in 1 % of the interface Li/Sc atoms in the training structures. 95 % bootstrap pivotal confidence intervals 65 were calculated using 1,000 bootstrap samples with 1,500 structures in each sample.
@@ 4.3 DFT calculations
All DFT calculations were performed using the Vienna Ab initio Simulation Package (VASP) 66, 67, 68, 69, with the projector augmented-wave (PAW) method 70, 71. The Li (), Sc (), Co (), Cl (), O (), Mg (), and S () electrons were treated as valence electrons in the pseudopotentials. The generalized gradient approximation (GGA) with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional72 was used. The orbitals were expanded using a plane wave basis with cutoff energy of 520 eV for bulk structures, and 450 eV for interface structures (to lower computational cost). The DFT+U approach 73 was used to account for the electron localization of the Co- states, and we selected a value of 4 eV as reported in other literature 74.
To investigate the origin of the interfacial disorder observed for chlorides, interfaces of simpler materials (LiO-Li, LiO-LiCl, and LiO-MgS) were studied. Structural relaxations were first performed on unit cells of LiO (), Li (), LiCl (), and MgS (). The cell shapes, cell volumes, and atom positions were allowed to relax, until the force on each atom was below 0.001 eV/Å. The Brillouin zone was sampled using a (8×8×8) Monkhorst-Pack k-point grid for Li, and a (6×6×6) Monkhorst-Pack k-point grid for LiO, LiCl, and MgS. The interface structures were constructed using the MPInterfaces package 75, which implements the lattice matching algorithm proposed by Zur et al. 76. The configurations of interfaces constructed and their lattice mismatches are listed in ESI Table S1. All interfaces were constructed with vacuum spacings of at least 14 Å. We chose the (100) orientation as a representative plane for Li, LiCl, and MgS.
To study the effect of different terminations of LiO surfaces on the interfacial structure with LiCl, we chose the LiO(100) orientation which can either be Li-terminated of O-terminated. The surfaces of the LiO(100) slabs are polar, so half of the Li/O atoms were moved from one surface to the other to neutralize the polarity (resulting in ‘Tasker Type 2b’ surfaces 77, 78). Structural relaxations were performed on LiO(100)-LiCl(100) interfaces for both terminations, allowing the cell shapes, cell volumes, and atom positions to relax until the force on each atom was below 0.05 eV/Å. The Brillouin zone was sampled using a (3×2×1) gamma-centered k-point grid. To study the effect of mechanical stiffness on interfacial structure, we chose the LiO(110) orientation, which exposes both Li and O at its surface, and paired it with Li(100), LiCl(100), and MgS(100). Structural relaxations were performed on these interfaces using the same procedure. The Brillouin zone was sampled using a (8×2×1), (3×2×1), and (3×2×1) gamma-centered k-point grid for LiO(110)-Li(100), LiO(110)-LiCl(100), and LiO(110)-MgS(100) respectively.
@@
For the LCO-LSC interface system, we began by performing structural relaxations on unit cells of LCO () and LSC ( 35), allowing the cell shapes, cell volumes, and atom positions to relax, until the force on each atom was below 0.001 eV/Å. The Brillouin zone was sampled using a (9×9×1) gamma-centered k-point grid for LCO, and a (6x4x6) Monkhorst-Pack k-point grid for LSC. To construct the interface structures, we chose the orientations of LCO(110), a Li fast-diffusing plane 79, and LSC(100), a representative plane. The surfaces of the LSC(100) slab are polar, so half of the Cl atoms were moved from one surface to the other to neutralize the polarity.
The configuration of the lattice-matched interface is given in ESI Table S1, where the average lattice mismatch is 2.17 %. In a preliminary calculation, the interface was constructed by simply placing the LCO(110) and LSC(100) slabs in contact with each other, and we obtained a relaxed structure similar to that depicted in Fig. 1b. An interface region with disordered LSC atoms formed with a thickness of around 5 Å. Using this structure as reference, we then generated random interface structures using the CALYPSO package 80, 81, 82. We used LCO(110) and LSC(100) slabs with 4 and 7 layers respectively. For each structure, an interface region thickness was randomly chosen between 4 and 6 Å, then the region was randomly populated with 3 formula units of LSC (9 Li, 3 Sc, and 18 Cl atoms). A random lateral displacement (in-plane direction parallel to the interface) was also applied to the LSC slab. A vacuum spacing of 14 Å was included in all the interface structures. For a subset of the generated structures, random mutual exchanges were performed between atoms in the interface region and the top layer of LCO, where we allowed up to 3 Sc Co, 3 Li Co, and 6 Cl O exchanges per structure. All interface structures have a total of 264 atoms per structure.
@@
A total of 350 structures were generated without mutual exchanges, and 600 structures were generated with mutual exchanges. Structural relaxations were performed on all interface structures, where the lateral lattice vectors were fixed to the optimized values of the LCO slab (the elastic moduli of LCO 83 is significantly larger than LSC 84). The relaxations were performed in two stages. In the first stage, a kinetic energy cutoff of 374.3 eV was used, and only the atoms in the interface region were allowed to relax for 50 ionic steps. In the second stage, a kinetic energy cutoff of 450 eV was used, and all atoms were allowed to relax until the force on each atom was below 0.1 eV/Å. The Brillouin zone was sampled at the gamma point only for both stages.
@@ 4.4 M3GNet
We trained the M3GNet interatomic potential 48 on the LCO-LSC interface structures to allow us to perform relaxations quickly. The dataset used for M3GNet training and evaluation included the 950 relaxed interface structures and 14,534 intermediate ionic steps from the DFT relaxations (second stage only). The intermediate steps were sampled starting from the 5th ionic step of every relaxation with an interval of 10 steps. The dataset was split into 80 % training, 10 % validation, and 10% test data. We used the Adam optimizer 85 to optimize the loss function, , as follows:
where MSE, MSE, and MSE are the mean squared error of energy, force and stress respectively. We tested different learning rates and batch sizes, and trained M3GNet models for 24 hours each using Tesla V100 GPUs. We found that the losses plateaued within the given training time, and there was little difference in errors between the different hyperparameters (see ESI Table S2). We chose the model with the smallest loss (learning rate = 0.001, batch size = 4) as the working interatomic potential of this work (see ESI Fig. S15 for the loss curves). The test set mean absolute errors for energy, force, and stress are 2.70 mev/atom, 20.9 meV/Å, and 0.0146 GPa respectively.
The GAN training dataset was obtained by randomly generating interface structures, then pre-relaxing them using the M3GNet interatomic potential. No mutual ion exchanges were performed when the structures were generated since the exchanges mostly led to high-interface-energy structures. The structures were optimized using a conjugate gradient algorithm 86 through the atomic simulation environment (ASE) package 87. The relaxations were stopped when the force on each atom was below 0.05 eV/Å. Then, the structures were relaxed through DFT calculations using a kinetic energy cutoff of 450 eV. Only the atom positions were allowed to relax until the force on each atom was below 0.1 eV/Å. The Brillouin zone was sampled at the gamma point only. 1,500 low-interface-energy structures were filtered from the relaxed structures as the training dataset. The initial 950 structures that were relaxed using DFT calculations only were just used for training the M3GNet interatomic potential, but not for GAN training. This is to ensure a fair comparison between the GAN-generated and training structures, since we intended to pre-relax the GAN-generated structures with the M3GNet interatomic potential. Table 3 shows the errors between the M3GNet- and DFT-calculated total energies of the interface structures. The errors are generally low even when compared to the final DFT-relaxed structures, showing that the M3GNet relaxations are able to give accurate predictions of the energy, and yield structures close to DFT convergence.
@@
mean |
std |
mean |
std |
mean |
std |
| low energy |
17.6 |
25.7 |
3.14 |
3.41 |
2.15 |
1.42 |
| high energy |
56.1 |
51.8 |
11.0 |
15.3 |
4.75 |
4.51 |
| all |
31.7 |
41.8 |
6.03 |
10.4 |
3.10 |
3.21 |
Before relaxing the GAN-generated structures, the structures were first preprocessed by moving apart atoms that were too close together until all atomic distances were 1.5 Å. For model comparison between architectures, the structures were relaxed using the M3GNet interatomic potential only (no DFT relaxation). For the best GAN model, the structures were relaxed using the M3GNet interatomic potential followed by DFT calculations. The normalized interface energies, , of relaxed structures were calculated using the following equation:
where is the total energy of the interface in eV, is the total number of atoms, and is the interface area. The interface energies are normalized such that structures with interface energy 0 J/m are considered to be low-interface-energy structures. The normalization threshold of -4.78 eV/atom was chosen such that the low-interface-energy structures fall within a range of 0.4 J/m. For reference, the observation frequency of a grain boundary in aluminum metal decreases by 95 % when the grain boundary energy increases by around 0.35 J/m 50.
@@
The CryinGAN code and datasets used in this study are available at https://github.com/ertekin-research-group/CryinGAN.
@@
A.X.B.Y. and E.E conceived the idea. A.X.B.Y. wrote the code, carried out the calculations/experiments, and performed the analysis. E.E. supervised and guided the project. The manuscript was prepared by A.X.B.Y. Both authors reviewed and edited the manuscript.

### v2

@@ 4.2.5 CryinGAN
The CryinGAN architecture is as shown in Fig. 6. CryinGAN was developed using a different interface dataset from the Dismai-Bench interface datatset, consisting of 1,500 low-interface-energy structures relaxed using M3GNet followed by DFT calculations. This is to facilitate evaluation of the best performing model through comparisons between DFT-relaxed training and generated structures. Benchmarking of CryinGAN was still carried out by training models on the Dismai-Bench datasets. During model development, CryinGAN models were trained with different values to study the effect of on model performance. The models were trained with a batch size of 32 for 100,000 epochs. Adam optimizers with learning rate of were used for the generator and discriminator(s). The generator was trained only once every 5 batches to help stabilize the training. CryinGAN-comb, CryinGAN-max, and CryinGAN-mix models were trained using the same procedure. CryinGAN-comb uses only a single combined discriminator, so no testing was required. CryinGAN-max and CryinGAN-mix were trained using = 0 and 0.05. For = 0, max pooling and mix pooling were used for the fractional coordinate discriminator of CryinGAN-max and CryinGAN-mix respectively. For = 0.05, max pooling and mix pooling were used for the bond distance discriminator of CryinGAN-max and CryinGAN-mix respectively, whereas average pooling was used for the fractional coordinate discriminator of both models. The various GAN models were compared using 1,000 structures generated from each model.
@@
We also tested a graph convolutional neural network architecture by adapting CGCNN 59 as the discriminator. Instead of predicting material properties, the output of CGCNN was used to estimate the Wasserstein distance between the training and generated structures. Default values for the hyperparameters were used, as provided in the code repository of CGCNN. The GAN was optimized using the Adam optimizer and a batch size of 32, where we tested learning rates of , , , and . We found that the losses diverged and the GAN was unable to generate meaningful structures. We did not further pursue the development of graph convolutional neural networks as discriminators, considering similar failures reported in literature for point cloud GANs 58.
The best GAN architecture was found to be CryinGAN using both discriminators. For the disordered interface, was found to be optimum (see ESI Fig. S15). For the alloys, was found to be optimum (see ESI Fig. S25). For amorphous Si, CryinGAN was unable to generate meaningful structures (see Fig. 3), so it was not benchmarked for amorphous Si. Benchmark models were trained on the Dismai-Bench disordered interface and alloy datasets using = 0.05 and 0.1 respectively. For the disordered interface dataset, we did not perform any data augmentation. CryinGAN is permutationally invariant, and the model does not need to learn to generate rotated interface structures. Each structure in the dataset was originally generated with the LCO slab fixed and the LSC slab randomly displaced laterally (parallel to the interface). Training models on a dataset with lateral translation augmentations was found to slow down CryinGAN’s learning, since it had to learn to generate structures with displaced LCO (and LSC) slabs. Therefore, we did not apply any translation augmentation as well. For the alloy datasets, we performed data augmentation by translating the structures by integer multiples of the unit cell lattice constant (3.6 Å). Each structure is a 444 supercell, giving unique translations, so we augmented each structure 63 times. The disordered interface and alloy benchmark models were trained for 100,000 and 1,500 epochs respectively, and models from the last epoch were used for benchmarking. All models were trained with a batch size of 32.
4.3 Dismai-Bench benchmarking
Benchmarking was performed by training all generative models on each dataset separately from scratch, such that the models only generate one type of structure at a time. For each generative model architecture, three separate models were trained for each dataset, and the benchmark metrics were averaged. 1,000 structures were generated for each model to calculate the benchmark metrics.
4.3.1 Disordered LSC-LCO interface
The generated interface structures were first post-processed by moving apart atoms that were too close together using an iterative algorithm. In each iteration, the algorithm determines all unique pairs of atoms too close together, and increases the magnitude of their bond vectors. The algorithm repeats itself until all atomic distances are 1.5 Å. We set a maximum of 100 iterations, and any structure that still had atoms too close together after 100 iterations was rejected. The structures were then relaxed using the M3GNet interatomic potential, allowing only the atom positions to relax. The relaxations were stopped when the force on each atom was below 0.05 eV/Å. The percentage of failed structures was calculated.
For each successfully relaxed structure, the CrystalNNFingerprint 55 was calculated for the cations (Li, Co, and Sc), allowing only the anions (Cl and O) to be considered as neighbors. We also appended the fraction of Cl and O neighbors in each motif to the fingerprints, so that the fingerprints contain both chemical and coordination information. Average fingerprints were calculated by averaging the site fingerprints of each structure, then averaging across all structures. The Euclidean distances between the average fingerprint of the generated structures and the training structures were calculated.
4.3.2 Amorphous Si
The generated amorphous Si structures were post-processed to move apart atoms too close together, using the same procedure described for the disordered interfaces. The structures were then relaxed using the SOAP-GAP interatomic potential, allowing only the atom positions to relax. The relaxations were stopped when the force on each atom was below 0.05 eV/Å. The percentage of failed structures was calculated.
For each successfully relaxed structure, the CrystalNNFingerprint 55 was calculated for all Si atoms. The Euclidean distances between the average fingerprint of the generated structures and the training structures were calculated. The RDF of each structure was calculated using the vasppy 85 package. The RDFs were calculated between 0.0 Å and 10.0 Å using a bin width of 0.02 Å. The RDFs were averaged across structures, and the Euclidean distances between the average RDF of the generated structures and the training structures were calculated. The bond angles of each structure was also calculated. The neighbors of each atom were determined using the CrystalNN 55 algorithm. The bond angle distribution was calculated by binning the bond angles using a bin width of 0.5∘. The bond angle distributions were averaged across structures, and the Euclidean distances between the average bond angle distribution of the generated structures and the training structures were calculated.
4.3.3 Disordered stainless steel alloy
The generated alloy structures were post-processed to remove atoms not on lattice sites. An atom was considered to be on a lattice site if it was within a 0.8 Å radius of the lattice site. If a lattice site already had an atom assigned to it, no other atom can be assigned to the site. Any structure with > 50 atoms removed (20 % of all atoms) was considered a failed structure and rejected. The percentages of failed structures and accepted structures with site vacancies were calculated.
The clusters (1 monomer and 7 dimers) of each structure were counted, where we considered up to the 7 neighbor shell. A fingerprint was constructed for each structure using the vector of conditional probabilities of observing each cluster (e.g., the probability of observing a 1st nearest neighbor Cr-Fe dimer given two nearest neighbor sites). The Euclidean distances between the average cluster fingerprint of the generated structures and the training structures were calculated.
4.4 Detailed evaluation of CryinGAN-generated interfaces
@@
We further evaluated the disordered interface structures generated by CryinGAN. We trained a CryinGAN model on the dataset of DFT-relaxed interface structures with . The generator was trained once every 2 batches to speed up the training. We found the training to be stable at this frequency as shown by the Wasserstein distance plot in ESI Fig. S26a. Although the Wasserstein distance plateaued early in the training, the percentage of (relaxed) low-interface-energy structures continued to gradually increase (see ESI Fig. S26b and c). We stopped the training when the energy improvements have mostly plateaued. Structures were generated using the model from the last epoch, and relaxed using M3GNet followed by DFT calculations. Low-interface-energy structures were filtered from the relaxed structures.
@@
Three datasets were used to perform the structural analysis: (1) CryinGAN-generated structures with low interface energy ( J/m2), (2) randomly generated structures with low interface energy (i.e., the training structures), and (3) randomly generated structures with high interface energy ( J/m2). Each dataset contains 1,500 structures that were relaxed using M3GNet followed by DFT calculations. The coordination environment of atoms in the interface region was compared between the datasets. The coordination motif fingerprints of Li and Sc atoms in the interface region were calculated for each dataset and averaged. The Euclidean distance and cosine similarity between the average fingerprint of the CryinGAN/high energy dataset and the training dataset were calculated. To visualize the coordination motif distribution of the datasets, the most likely coordination motif of each interface Li/Sc atom was identified by first selecting the coordination number with the highest likelihood, then selecting the coordination motif of this coordination number with the highest local structure order parameter (LoStOP) 55. We only plotted coordination motifs that appeared in 1 % of the interface Li/Sc atoms in the training structures. 95 % bootstrap pivotal confidence intervals 86 were calculated using 1,000 bootstrap samples with 1,500 structures in each sample.
@@ 4.5 DFT calculations
All DFT calculations were performed using the Vienna Ab initio Simulation Package (VASP) 87, 88, 89, 90, with the projector augmented-wave (PAW) method 91, 92. The Li (), Sc (), Co (), Cl (), and O () electrons were treated as valence electrons in the pseudopotentials. The generalized gradient approximation (GGA) with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional93 was used. The orbitals were expanded using a plane wave basis with cutoff energy of 520 eV for bulk structures, and 450 eV for interface structures (to lower computational cost). The DFT+ approach 94 was used to account for the electron localization of the Co- states, and we selected a value of 4 eV as reported in other literature 95.
@@
We first performed structural relaxations on unit cells of LCO () and LSC ( 41), allowing the cell shapes, cell volumes, and atom positions to relax, until the force on each atom was below 0.001 eV/Å. The Brillouin zone was sampled using a (9×9×1) gamma-centered k-point grid for LCO, and a (6×4×6) Monkhorst-Pack k-point grid for LSC. The relaxed unit cells were used to construct the LCO(110) and LSC(100) slabs as described in Section 4.1.3.
@@
To generate the dataset for training the M3GNet interatomic potential, LSC(100)-LCO(110) interfaces were randomly generated as described in Section 4.1.3. A total of 350 structures were generated without mutual exchanges, and 600 structures were generated with mutual exchanges. Random mutual exchanges were performed between atoms in the interface region and the top layer of LCO, where we allowed up to 3 Sc Co, 3 Li Co, and 6 Cl O exchanges per structure. Structural relaxations were performed on all interface structures with the cell shapes and cell volumes fixed. The relaxations were performed in two stages. In the first stage, a kinetic energy cutoff of 374.3 eV was used, and only the atoms in the interface region were allowed to relax for 50 ionic steps. In the second stage, a kinetic energy cutoff of 450 eV was used, and all atoms were allowed to relax until the force on each atom was below 0.1 eV/Å. The Brillouin zone was sampled at the gamma point only for both stages.
For structures that have been pre-relaxed using M3GNet (after the M3GNet model was trained), DFT relaxations were performed without the first stage relaxation (second stage only). The interface dataset for Dismai-Bench was created by relaxing randomly generated structures with M3GNet only. The interface dataset for developing CryinGAN was created by relaxing randomly generated structures with M3GNet followed by DFT calculations. No mutual ion exchanges were performed when the structures were generated since the exchanges mostly led to high-interface-energy structures. The initial 950 structures that were relaxed using DFT calculations only were just used for training the M3GNet model, but not any generative model.
@@ 4.6 M3GNet
We trained the M3GNet interatomic potential 49 on the LSC-LCO interface structures to allow us to perform relaxations quickly. The dataset used for M3GNet training and evaluation included the 950 relaxed interface structures and 14,534 intermediate ionic steps from the DFT relaxations (second stage only). The intermediate steps were sampled starting from the 5th ionic step of every relaxation with an interval of 10 steps. The dataset was split into 80 % training, 10 % validation, and 10 % test data. We used the Adam optimizer to optimize the loss function, , as follows:
where MSE, MSE, and MSE are the mean squared error of energy, force and stress respectively. We tested different learning rates and batch sizes, and trained M3GNet models for 24 hours each. We found that the losses plateaued within the given training time, and there was little difference in errors between the different hyperparameters (see ESI Table S6). We chose the model with the smallest loss (learning rate = 0.001, batch size = 4) as the working interatomic potential of this work (see ESI Fig. S27 for the loss curves). The test set mean absolute errors for energy, force, and stress are 2.70 meV/atom, 20.9 meV/Å, and 0.0146 GPa respectively.
We compared interface structures that were pre-relaxed with M3GNet with their final structures after subsequent DFT relaxation. Table 5 shows the errors between the M3GNet- and DFT-calculated total energies. The errors are generally low even when compared to the final DFT-relaxed structures, showing that the M3GNet relaxations are able to give accurate predictions of the energy, and yield structures close to DFT convergence.
@@
| mean | std | mean | std | mean | std | |
| low energy | 17.6 | 25.7 | 3.14 | 3.41 | 2.15 | 1.42 |
| high energy | 56.1 | 51.8 | 11.0 | 15.3 | 4.75 | 4.51 |
| all | 31.7 | 41.8 | 6.03 | 10.4 | 3.10 | 3.21 |
@@
The datasets and interatomic potentials used are available openly at https://doi.org/10.5281/zenodo.12710372. The Dismai-Bench benchmarking and generative model code used in this work are available at https://github.com/ertekin-research-group/Dismai-Bench. The CryinGAN code is available separately at https://github.com/ertekin-research-group/CryinGAN.
@@
A.X.B.Y. and E.E conceived the idea. A.X.B.Y. designed the benchmark, wrote the code, carried out the calculations/experiments, and performed the analysis. T.S. assisted with the alloy dataset preparation and code for cluster counting. E.E. supervised and guided the project. The manuscript was prepared by A.X.B.Y. and T.S. All authors reviewed and edited the manuscript.